if ceasy is 6 years older than her friend Lisa, and john is 5 years older than ceasy,and total of their age is 41,Then what is the Lisa age?
Answer by TimothyLamb(4379)

---
x = ceasy
y = lisa
z = john
---
x = y + 6
z = x + 5
x + y + z = 41
---
put the system of linear equations into standard form
---
x - y = 6
x - z = -5
x + y + z = 41
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = ceasy = 14
y = lisa = 8
z = john = 19
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php